day shift, 7 July 99

Marj Corcoran (corcoran@physics.rice.edu)
Wed, 7 Jul 1999 15:31:06 -0500 (CDT)


Shift Summary for day shift, 07/07/99

On Shift: J. LaDue, M. Corcoran, C. Wolfe

Hours of beam available at all intensities: 7.5
Hours of beam available at nominal intensity: 7.0
Hours of E832 data (special runs not included): 7.5
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: Beam off when TeV ramp lost; linac RF
problems

========================================================================
Runs ended this shift:
Run 13834 E832 data in progress

Tapes brought to FCC this shift: None
========================================================================
Accesses during shift: None
========================================================================

Overview of shift: Very quiet shift, beam stable during most of shift.
However, BOA was around 1.4 to 1.5 all day, even though we called MCR about
it several times. SEM was 4.5 to 5.5 E12 during most of the shift.

Gas flow to DC4 was increased slightly to see if it helps inefficient area
in plane 15.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13834 180 8:00 BOA remains high, has been around 1.5 this shift

13834 400 13:00 ONSPLIT requests tape mount QKCC04 on KTEV1

13834 415 13:30 Trigger error 10, KFREND cures

13834 447 14:10 TeV lost ramp; est 5 mins

13834 450 14:15 Beam is back, but very low intensity. Returns to
nominal intensity in about 10 mins.

13834 485 15:05 Linac RF problems, beam gone, no estimate
15:15 Estimate 30 mins for beam

========================================================================
Follow up Notes on Previous Shift
========================================================================
Items to Follow up on Next Shift